OpenStudy (anonymous):

With terms of 2/10, n/30 means

OpenStudy (anonymous):

6/30

OpenStudy (anonymous):

Could you please clarify your question because it's very ambiguous... Did you mean to say: With a term of n=2/10=0.2 what does n/30 mean? Or maybe, with either n=2 or n=10, what does n/30 mean? There are a few other possible interpretations... =/

OpenStudy (anonymous):

2/10=n/30 =>n=6 I've understood the task in such a way....

OpenStudy (anonymous):

but English isn't my native language, so I could have interpreted the task in a wrong way(

OpenStudy (anonymous):

Well did you already type out exactly what the question says?

OpenStudy (anonymous):

What did the question say exactly?